OpenStudy (anonymous):

If ax*x + bx +c = 0, then what is x?

OpenStudy (anonymous):

it is a standard equation there is no value of x

OpenStudy (anonymous):

(a)x*x + (b)x + (c) = 0 d*d = b*b - 4ac (-b +/- d)/2a=x

hero (hero):

x = quadratic formula

OpenStudy (anonymous):

\[x = \frac{\sqrt{b^2-4ac}-b}{2a}\]or\[x=\frac{-\sqrt{b^2-4ac}-b}{2a}\]

OpenStudy (anonymous):

can use quadratic formula

OpenStudy (anonymous):

the above formula is wrong

OpenStudy (anonymous):

\[x=\sqrt{b ^{2}-4ac}\div2a\]

hero (hero):

It's not wrong, it's just written in a strange way. Princess you need the +/- b as well

OpenStudy (anonymous):

yess ofcourse +,- as well